Prepare journal entries to record the following transactions entered into by the Novak Corp.: (Credit account titles are auto2022
Answer #1
June 1 Notes Receivable 15000
     Accounts Receivable—D. Gore 15000
Nov. 1 Accounts Receivable—Barlow, Inc.   17000
      Sales 17000
Nov. 5 Sales Returns and Allowances 1000
       Accounts Receivable—Barlow, Inc. 1000
Nov. 9 Cash 15520
Sales Discounts 480 =16000*3%
       Accounts Receivable—Barlow, Inc. 16000
Dec 31 Interest Receivable 525 =15000*6%*7/12
      Interest Revenue 525
2022
June 1 Cash 15900
      Interest Receivable 525
      Interest Revenue 375 =15000*6%*5/12
      Notes Receivable 15000